# Basic representations of genus zero nonabelian Hodge spaces

### v2

@@ 3.1. Generic form
If $I=\langle q\rangle_{\infty}$ is of generic form, the Fourier sphere coefficient $\lambda(I)$ of $I$ is the unique complex number $a\in\mathbb{C}$ such that $\langle\frac{a}{2}z^{2}+q\rangle$ has slope $<2$. Explicitly, this means that $q$ is of the form
$q=-\frac{a}{2}+q_{<2},$
where the exponential factor $q_{<2}$ has slope $<2$. We will denote $I_{<2}:=\langle q_{<2}\rangle$.
@@
For $A=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{SL}_{2}(\mathbb{C})$ let $h_{A}:\mathbb{P}^{1}\to\mathbb{P}^{1}$ denote the corresponding homography
$h_{A}:z\mapsto\frac{az+b}{cz+d}.$
@@
The following result says that $\operatorname{SL}_{2}(\mathbb{C})$ acts on the Fourier sphere by homographies:
@@
Let $I\in\pi_{0}(\mathcal{I})$ be a Stokes circle, $a=\lambda(I)$ its Fourier sphere coefficient, and $A\in SL_{2}(\mathbb{C})$. Then
$\lambda(A\cdot I)=h_{A}(\lambda(I)).$
@@
In the case where $a\neq\infty$ and $h_{A}(a)\neq\infty$, this is the result of Proposition 4.13 of [18], and one checks that this is still true in the remaining cases. ∎
This immediately implies the following:
@@
In particular a connection of generic form only has a singularity at infinity, i.e. $\bm{\breve{\Theta}}=\Theta_{\infty}$. A global modified irregular class $\bm{\breve{\Theta}}$ is of generic form if and only if $\lambda(\bm{\breve{\Theta}})\subset\mathbb{C}=\mathbb{P}^{1}\smallsetminus\{\infty\}$, where $\lambda(\bm{\breve{\Theta}})$ denotes the set of Fourier sphere coefficients of all its Stokes circles. This immediately implies:
@@
Let $\bm{\breve{\Theta}}$ be an effective modified irregular class. Then for $A$ in an open dense subset of $\operatorname{SL}_{2}(\mathbb{C})$, $A\cdot\bm{\breve{\Theta}}$ is of generic form.
@@
Let $(\Theta,\mathcal{C})$ be formal data at infinity with $\Theta$ of generic form, and such that $\lambda(\Theta)=\{a\}$ for some $a\in\mathbb{C}$. Let $A\in\operatorname{SL}_{2}(\mathbb{C})$. Let $\Theta^{<2}:=T_{-a}\cdot\Theta$ be the slope $<2$ irregular class obtained from $\Theta$ by subtracting $-\frac{a}{2}z^{2}$ to each of its Stokes circles. We have:
@@
    If $h_{A}(a)=\infty$, then $A\cdot(\Theta,\mathcal{C})$ is an admissible deformation of $F\cdot(\Theta^{<2},\mathcal{C})$.
@@
    Otherwise, if $h_{A}(a)\neq\infty$, then $A\cdot(\Theta,\mathcal{C})$ is an admissible deformation of $(\Theta,\mathcal{C})$ , or equivalently of $(\Theta^{<2},\mathcal{C})$.
@@
This is basically a consequence of the fact that any element in $\operatorname{SL}_{2}(\mathbb{C})$ can be factorised as a product of elementary transformations, and among those only the Fourier transform may act nontrivially on global fission data. More precisely, we can use lemma 4.15 of [18] to write down explicit factorisations of $A$. If $h_{A}(a)=\infty$, then we have $h_{AT_{a}F^{-1}}(\infty)=\infty$, so from this lemma, there exists $v\in\mathbb{C}^{*},\rho\in\mathbb{C}$ such that $AT_{a}F^{-1}=S_{v}T_{\rho}$, so that $A=S_{v}T_{\rho}FT_{-a}$. This implies $A\cdot(\Theta,\mathcal{C})=S_{v}T_{\rho}FT_{-a}\cdot(\Theta,\mathcal{C})=S_{v}T_{\rho}F\cdot(\Theta^{<2},\mathcal{C})$, and since twists and scalings do not change the equivalence class of admissible deformations $A\cdot(\Theta,\mathcal{C})$ is an admissible deformation of $F\cdot(\Theta^{<2},\mathcal{C})$.
Let us now consider the case $h_{A}(a)\neq\infty$. If $h_{A}(\infty)\neq\infty$ then from the lemma $A$ admits a factorisation $A=S_{\nu}T_{\rho}$ with $v\in\mathbb{C}^{*},\rho\in\mathbb{C}$, which implies that $A\cdot\Theta$ is an admissible deformation of $\Theta$. Otherwise, from the lemma, $A$ admits a factorisation $A=T_{\mu}FS_{\nu}T_{\rho}$, with $v\in\mathbb{C}^{*},\mu,\rho\in\mathbb{C}$. In turn $\Theta^{\prime}:=S_{\nu}T_{\rho}\cdot\Theta$ is an admissible deformation of $\Theta$, and $\lambda(\Theta^{\prime})=\{a^{\prime}\}$, with $a^{\prime}=h_{S_{\nu}T_{\rho}}(a)$, and we have $a^{\prime}\neq 0$ otherwise we would have $h_{A}(a)=h_{S_{\nu}F}(a^{\prime})=\infty$. Now, applying the propositions 2.14 and 2.15 to $\Theta^{\prime}$, the fact that all exponential factors in $\Theta^{\prime}$ are of slope 2 and have the same nonzero leading coefficient implies that this is also the case for $F\cdot\Theta^{\prime}$, and that $F\cdot(\Theta^{\prime},\mathcal{C})$ is an admissible deformation of $(\Theta^{\prime},\mathcal{C})$. Since $T_{\mu}$ doesn’t change the equivalence class of admissible deformations, we finallly obtain that $A\cdot(\Theta,\mathcal{C})$ is an admissible deformation of $(\Theta,\mathcal{C})$, which concludes the proof. ∎
In particular, we can write down explicit formulas for the rank at infinity of the elements of the orbit of $\Theta$. Let $N$ be the set of Stokes circles of $\Theta$, and let us set:
$N^{+}:=\{\langle q\rangle\in N\;|\;\operatorname{slope}(q+\frac{a}{2}z^{2})>1\}\subset N,\qquad N^{-}:=\{\langle q\rangle\in N\;|\;\operatorname{slope}(q+\frac{a}{2}z^{2})\leq 1\}\subset N,$
so that $N=N^{+}\sqcup N^{-}$. Let us write the elements of $N^{-}$ as
$I^{-}_{j}=\langle-\frac{a}{2}z^{2}+b_{j}z+q^{-}_{j}\rangle$
for $j=1,\dots,|N^{-}|$, where $\operatorname{slope}(q^{-}_{j})<1$, and set $s^{-}_{j}:=\operatorname{Irr}(q^{-}_{j})$, $r^{-}_{j}:=\operatorname{ram}(q^{-}_{j})$. Similarly, we write
$I^{+}_{j}=\langle-\frac{a}{2}z^{2}+q^{+}_{j}\rangle,$
for $j=1,\dots,|N^{+}|$, with $\operatorname{slope}(q^{+}_{j})>1$, and set $s^{+}_{j}:=\operatorname{Irr}(q^{+}_{j})$, $r^{+}_{j}:=\operatorname{ram}(q^{+}_{j})$.
@@
Let $n^{\pm}_{j}$ denote the multiplicities of the circles in $N^{\pm}$, and set $R^{+}_{j}:=s^{+}_{j}-r^{+}_{j}$, for $j=1,\dots,|N^{+}|$. Let $A\in\operatorname{SL}_{2}(\mathbb{C})$. We have
@@
This follows directly from the previous lemma and proposition 2.14, by observing that for the case $h_{A}(a)=\infty$, the Stokes circles in $N^{-}$ are sent by $A$ to finite distance (with the position of $A\cdot I_{j}^{-}$ determined by the coefficient $b_{j}$), while the circles in $N^{+}$ become circles of slope >1 at infinity, with $A\cdot I_{j}^{+}$ having ramification order $R_{j}^{+}$. ∎
3.2. The classes of nearby representations
We now discuss general orbits. Let $(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}})$ be global modified formal data, with $\bm{\breve{\Theta}}=\Theta_{\infty}$ be an effective irregular class of generic form. Let $a_{1},\dots a_{k}\in\mathbb{C}$ be the distinct Fourier sphere coefficients of its Stokes circles, i.e. we have $\lambda(\Theta_{\infty})=\{a_{1},\dots,a_{k}\}$.
Let $N$ be the set of Stokes circles of $\bm{\breve{\Theta}}$. For $I\in N$ we denote by $N_{I}\in\mathbb{N}_{>0}$ the multiplicity of $I$ in $\bm{\breve{\Theta}}$. We partition $N$ as follows: for $i=1,\dots,k$, we set
@@ 3.2. The classes of nearby representations
    $N_{i}^{+}:=\{\langle q\rangle\in N_{i}\;|\;\operatorname{slope}(q+\frac{a_{i}}{2}z^{2})>1\}\subset N_{i}$,
@@
    $N_{i}^{-}:=\{\langle q\rangle\in N_{i}\;|\;\operatorname{slope}(q+\frac{a_{i}}{2}z^{2})\leq 1\}\subset N_{i}$,
so that
$N=N_{1}\sqcup\dots\sqcup N_{k},\qquad N_{i}=N_{i}^{+}\sqcup N_{i}^{-}.$
We also set $N^{+}:=N_{1}^{+}\sqcup\dots\sqcup N_{k}^{+}$, and $N^{-}:=N_{1}^{-}\sqcup\dots\sqcup N_{k}^{-}$. For $i=1,\dots,k$, let $\Theta_{i}:=\sum_{I\in N_{i}}n_{I}I$ be the irregular class obtained from $\Theta_{\infty}$ obtained by discarding all circles with Fourier sphere coefficient different from $a_{i}$, $\mathcal{C}_{i}$ the corresponding collection of conjugacy classes, and let $\Theta^{<2}_{i}$ be the irregular class at infinity obtained from $\Theta_{i}$ by subtracting $-\frac{a_{i}}{2}z^{2}$ to each Stokes circle (in particular all circles of $\Theta_{i}$ have slope $<2$).
@@
    If $h_{A}(a_{i})=\infty$ for some $i\in\{1,\dots,k\}$, then $A\cdot(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}})$ is an admissible deformation of
    $F\cdot(\Theta_{i}^{<2},\mathcal{C}_{i})+\sum_{l\neq i}(\Theta_{l},\mathcal{C}_{l}).$
@@
    Otherwise, if $h_{A}(a_{i})\neq\infty$ for each $i=1,\dots,k$, then $A\cdot(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}})$ is an admissible deformation of $(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}})$.
@@
This motivates the following definition:
@@
Let $\mathbf{T}$ be a fission tree of generic form, $N$ its set of leaves and $\bm{\mathcal{C}}$ the datum of a conjugacy class $\mathcal{C}_{i}\in\operatorname{GL}_{n_{i}}(\mathbb{C})$ for each leaf $i\in N$. Let $\Theta_{\infty}$ be an irregular class at infinity of generic form with fission tree $\mathbf{T}$. Keeping previous notations, for $i=1,\dots,k$, let $(\mathbf{F}_{i},\bm{\mathcal{C}}_{i})$ be the class of admissible deformations of the nonmodified formal data corresponding to the formal data
$F\cdot(\Theta_{i}^{<2},\mathcal{C})+\sum_{l\neq i}(\Theta_{l},\mathcal{C}).$
The definition makes sense, indeed it follows from Propositions 2.14 and 2.15 that the class of admissible deformations of $F\cdot(\Theta_{i}^{<2},\mathcal{C})+\sum_{l\neq i}(\Theta_{l},\mathcal{C})$ only depends on the fission tree $\mathbf{T}$. We call $(\mathbf{F}_{i},\bm{\mathcal{C}}_{i})$ the $i$-th nongeneric reading of $\mathcal{(}\mathbf{T},\bm{\mathcal{C}})$. The previous theorem can then be reformulated as follows:
@@
Let $(\mathbf{\Theta},\bm{\mathcal{C}})$ be a wild Riemann surface with boundary data on $\mathbb{P}^{1}$ of generic form, with class of admissible deformations ($(\{\mathbf{T}\},\bm{\mathcal{C}})$. Then the classes of its nearby representations of $\mathcal{M}(\mathbf{\Theta},\bm{\mathcal{C}})$ are ($(\{\mathbf{T}\},\bm{\mathcal{C}})$ and $(\mathbf{F}_{i},\bm{\mathcal{C}}_{i})$ for $i=1,\dots k$.
By abuse of language, we will sometimes allow ourselves to say that $\{\mathbf{T}\}$ and $\mathbf{F}_{i}$, for $i=1,\dots,k$ are the nearby representations of $\mathbf{T}$.
@@
Let $(E,\nabla)$ be an arbitrary algebraic connection on a Zariski open subset of $\mathbb{P}^{1}$, with global modified irregular class $(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}})$, $A\in\operatorname{SL}_{2}(\mathbb{C})$ such that $A\cdot(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}})$ is of generic form. Let $\mathbf{T}$ be the (short) fission tree of $A\cdot\bm{\breve{\Theta}}$, and $\bm{\mathcal{C}}$ the collection of conjugacy classes of $A\cdot(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}})$. Then the pair $(\mathbf{T},\bm{\mathcal{C}})$ does not depend on the choice of $A$.
@@
This is an immediate consequence of Thm. 3.8. ∎
@@
Keeping the setup of the proposition, we say that $(\mathbf{,}\bm{\mathcal{C}})$ is the enriched tree of $(E,\nabla)$ and write $\mathscr{T}(E,\nabla):=(\mathbf{T},\bm{\mathcal{C}})$.
Prop. 3.11 implies that the enriched tree is invariant under symplectic transformations.
@@
If $(E,\nabla)$ is irreducible and is not a rank one connection with just one pole of order less than 2 at infinity, so that $A\cdot(E,\nabla)$ is well-defined for any $A\in\operatorname{SL}_{2}(\mathbb{C})$, we have
$\mathscr{T}(A\cdot(E,\nabla))=\mathscr{T}(E,\nabla).$
3.3. Properties of classes of nearby representations
We can provide explicit formulas for the rank of the nearby representations. For $i=1,\dots,k$, let us denote by $\langle q^{+}_{i,j}\rangle$ the elements of $N_{i}^{+}$, and write $\langle q^{+}_{i,j}\rangle=\langle-\frac{a_{i}}{2}z^{2}+q^{\prime+}_{ij}\rangle$, with $q^{\prime+}_{ij}$ having ramification $r^{+}_{ij}$ and slope $s^{+}_{ij}/r^{+}_{ij}>1$. Similarly, let us denote by $\langle q^{-}_{i,j}\rangle$ the elements of $N_{i}^{-}$, and write $\langle q^{-}_{i,j}\rangle=\langle-\frac{a_{i}}{2}z^{2}+b_{ij}z+q^{\prime-}_{ij}\rangle$, with $q^{\prime-}_{ij}$ having ramification $r^{+}_{ij}$ and slope $s^{+}_{ij}/r^{+}_{ij}<1$. Let $n_{ij}^{\pm}$ be the multiplicities of the active circles.

### v3

@@ 3.1 Generic form
If $I=\bigl\langle q\bigr\rangle_{\infty}$ is of generic form, the Fourier sphere coefficient $\lambda(I)$ of $I$ is the unique complex number $a\in\mathbb{C}$ such that $\bigl\langle\frac{a}{2}z^{2}+q\bigr\rangle$ has slope <2<2. Explicitly, this means that $q$ is of the form
$q=-\frac{a}{2}z^{2}+q_{<2},$
where the exponential factor $q_{<2}$ has slope <2<2. We will denote $I_{<2}:=\bigl\langle q_{<2}\bigr\rangle$.
@@
For $A=\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\in\operatorname{SL}_{2}(\mathbb{C})$, let $h_{A}\colon\mathbb{P}^{1}\to\mathbb{P}^{1}$ denote the corresponding homography
$h_{A}\colon\ z\mapsto\frac{az+b}{cz+d}.$
@@
The following result says that $\operatorname{SL}_{2}(\mathbb{C})$ acts on the Fourier sphere by homographies.
@@
Let $I\in\pi_{0}(\mathcal{I})$ be a Stokes circle, $a=\lambda(I)$ its Fourier sphere coefficient, and $A\in SL_{2}(\mathbb{C})$. Then $\lambda(A\cdot I)=h_{A}(\lambda(I))$.
@@
In the case where $a\neq\infty$ and $h_{A}(a)\neq\infty$, this is the result of [20, Proposition 4.13], and one checks that this is still true in the remaining cases. ∎
This immediately implies the following.
@@
In particular, a connection of generic form only has a singularity at infinity, i.e., $\bm{\breve{\Theta}}=\Theta_{\infty}$. A global modified irregular class $\bm{\breve{\Theta}}$ is of generic form if and only if $\lambda\bigl(\bm{\breve{\Theta}}\bigr)\subset\mathbb{C}=\mathbf{P}^{1}\smallsetminus\{\infty\}$, where $\lambda\bigl(\bm{\breve{\Theta}}\bigr)$ denotes the set of Fourier sphere coefficients of all its Stokes circles. This immediately implies the following.
@@
Let $\bm{\breve{\Theta}}$ be a global modified irregular class. Then for $A$ in an open dense subset of $\operatorname{SL}_{2}(\mathbb{C})$, $A\cdot\bm{\breve{\Theta}}$ is of generic form.
@@
Let $(\Theta,\mathcal{C})$ be formal data at infinity with $\Theta$ of generic form, and such that $\lambda(\Theta)=\{a\}$} for some $a\in\mathbb{C}$. Let $A\in\operatorname{SL}_{2}(\mathbb{C})$. Let $\Theta^{<2}:=T_{-a}\cdot\Theta$ be the slope <2<2 irregular class obtained from $\Theta$ by subtracting $\frac{a}{2}z^{2}$ to each of its Stokes circles. We have
@@
    If $h_{A}(a)=\infty$, then $A\cdot(\Theta,\mathcal{C})$ is an admissible deformation of $F\cdot\bigl(\Theta^{<2},\mathcal{C}\bigr)$.
@@
    Otherwise, if $h_{A}(a)\neq\infty$, then $A\cdot(\Theta,\mathcal{C})$ is an admissible deformation of $(\Theta,\mathcal{C})$ , or equivalently of $\bigl(\Theta^{<2},\mathcal{C}\bigr)$.
@@
This is basically a consequence of the fact that any element in $\operatorname{SL}_{2}(\mathbb{C})$ can be factorised as a product of elementary transformations, and among those only the Fourier transform may act nontrivially on global fission data. More precisely, we can use [20, Lemma 4.15] to write down explicit factorisations of $A$. If $h_{A}(a)=\infty$, then we have $h_{AT_{a}F^{-1}}(\infty)=\infty$, so from this lemma, there exists $v\in\mathbb{C}^{*},\rho\in\mathbb{C}$ such that $AT_{a}F^{-1}=S_{v}T_{\rho}$, so that $A=S_{v}T_{\rho}FT_{-a}$. This implies $A\cdot(\Theta,\mathcal{C})=S_{v}T_{\rho}FT_{-a}\cdot(\Theta,\mathcal{C})=S_{v}T_{\rho}F\cdot\bigl(\Theta^{<2},\mathcal{C}\bigr)$, and since twists and scalings do not change the equivalence class of admissible deformations $A\cdot(\Theta,\mathcal{C})$ is an admissible deformation of $F\cdot\bigl(\Theta^{<2},\mathcal{C}\bigr)$.
Let us now consider the case $h_{A}(a)\neq\infty$. If $h_{A}(\infty)\neq\infty$ then from the lemma $A$ admits a factorisation $A=S_{\nu}T_{\rho}$ with $v\in\mathbb{C}^{*},\rho\in\mathbb{C}$, which implies that $A\cdot\Theta$ is an admissible deformation of $\Theta$. Otherwise, from the lemma, $A$ admits a factorisation $A=T_{\mu}FS_{\nu}T_{\rho}$, with $v\in\mathbb{C}^{*}$}, $\mu,\rho\in\mathbb{C}$}. In turn $\Theta^{\prime}:=S_{\nu}T_{\rho}\cdot\Theta$ is an admissible deformation of $\Theta$, and $\lambda(\Theta^{\prime})=\{a^{\prime}\}$, with $a^{\prime}=h_{S_{\nu}T_{\rho}}(a)$}, and we have $a^{\prime}\neq 0$ otherwise we would have $h_{A}(a)=h_{S_{\nu}F}(a^{\prime})=\infty$. Now, applying Propositions 2.16 and 2.17 to $\Theta^{\prime}$, the fact that all exponential factors in $\Theta^{\prime}$ are of slope 2 and have the same nonzero leading coefficient implies that this is also the case for $F\cdot\Theta^{\prime}$, and that $F\cdot(\Theta^{\prime},\mathcal{C})$ is an admissible deformation of $(\Theta^{\prime},\mathcal{C})$. Since $T_{\mu}$ does not change the equivalence class of admissible deformations, we finally obtain that $A\cdot(\Theta,\mathcal{C})$ is an admissible deformation of $(\Theta,\mathcal{C})$, which concludes the proof. ∎
In particular, we can write down explicit formulas for the rank at infinity of the elements of the orbit of $\Theta$. Let $N$ be the set of Stokes circles of $\Theta$, and let us set
$N^{+}:=\left\{\bigl\langle q\bigr\rangle\in N\mid\operatorname{slope}\left(q+\frac{a}{2}z^{2}\right)>1\right\}\subset N,$
$N^{-}:=\left\{\bigl\langle q\bigr\rangle\in N\mid\operatorname{slope}\left(q+\frac{a}{2}z^{2}\right)\leq 1\right\}\subset N,$
[…]
@@
Let $n^{\pm}_{j}$ denote the multiplicities of the circles in $N^{\pm}$, and set $R^{+}_{j}:=s^{+}_{j}-r^{+}_{j}$, for $j\in\big\{1,\dots,\big|N^{+}\big|\big\}$. Let $A\in\operatorname{SL}_{2}(\mathbb{C})$. We have
@@
This follows directly from Lemma 3.5 and Proposition 2.16, by observing that for the case $h_{A}(a)=\infty$, the Stokes circles in $N^{-}$ are sent by AA to finite distance (with the location of $A\cdot I_{j}^{-}$ determined by the coefficient $b_{j}$), while the circles in $N^{+}$ become circles of slope >1>1 at infinity, with $A\cdot I_{j}^{+}$ having ramification order $R_{j}^{+}$. ∎
3.2 The classes of nearby representations
We now discuss general orbits. Let $\bigl(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}}\bigr)$ be effective global modified formal data, with $\bm{\breve{\Theta}}=\Theta_{\infty}$ of generic form. Let $a_{1},\dots,a_{k}\in\mathbb{C}$ be the distinct Fourier sphere coefficients of its Stokes circles, i.e., we have $\lambda(\Theta_{\infty})=\{a_{1},\dots,a_{k}\}\subset\mathbf{P}^{1}$.
Let $N$ be the set of Stokes circles of $\bm{\breve{\Theta}}$. For $I\in N$, we denote by $n_{I}\in\mathbb{N}_{>0}$ the multiplicity of $I$ in $\bm{\breve{\Theta}}$. We partition $N$ as follows: for $i\in\{1,\dots,k\}$, we set
@@ 3.2 The classes of nearby representations
    $N_{i}^{+}:=\big\{\bigl\langle q\bigr\rangle\in N_{i}\mid\operatorname{slope}\bigl(q+\frac{a_{i}}{2}z^{2}\bigr)>1\big\}\subset N_{i}$,
@@
    $N_{i}^{-}:=\big\{\bigl\langle q\bigr\rangle\in N_{i}\mid\operatorname{slope}\bigl(q+\frac{a_{i}}{2}z^{2}\bigr)\leq 1\big\}\subset N_{i}$,
so that $N=N_{1}\sqcup\dots\sqcup N_{k}$, $N_{i}=N_{i}^{+}\sqcup N_{i}^{-}$. We also set $N^{+}:=N_{1}^{+}\sqcup\dots\sqcup N_{k}^{+}$, and $N^{-}:=N_{1}^{-}\sqcup\cdots\allowbreak\sqcup N_{k}^{-}$. For $i\in\{1,\dots,k\}$, let $\Theta_{i}:=\sum_{I\in N_{i}}n_{I}I$ be the irregular class obtained from $\Theta_{\infty}$ obtained by discarding all circles with Fourier sphere coefficient different from $a_{i}$, $\mathcal{C}_{i}$ the corresponding collection of conjugacy classes, and let $\Theta^{<2}_{i}$ be the irregular class at infinity obtained from $\Theta_{i}$ by subtracting $\frac{a_{i}}{2}z^{2}$ to each Stokes circle (in particular all circles of $\Theta_{i}$ have slope <2<2).
@@
    If $h_{A}(a_{i})=\infty$ for some $i\in\{1,\dots,k\}$, then $A\cdot\bigl(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}}\bigr)$ is an admissible deformation of
    $F\cdot\bigl(\Theta_{i}^{<2},\mathcal{C}_{i}\bigr)+\sum_{l\neq i}(\Theta_{l},\mathcal{C}_{l}).$
@@
    Otherwise, if $h_{A}(a_{i})\neq\infty$ for each $i\in\{1,\dots,k\}$, then $A\cdot\bigl(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}}\bigr)$ is an admissible deformation of $\bigl(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}}\bigr)$.
@@
We notice that the corresponding classes of admissible deformations are in fact completely determined by the (short) fission tree of $\bm{\breve{\Theta}}$.
Proposition 3.9.
Keeping previous notations, let $\mathbf{T}$ be the fission tree of $\bm{\breve{\Theta}}$. Then for $i\in\{1,\dots,k\}$, the class of admissible deformations $(\mathbf{F}_{i},\bm{\mathcal{C}}_{i})$ of the nonmodified formal data associated to the formal data
$F\cdot\bigl(\Theta_{i}^{<2},\mathcal{C}_{i}\bigr)+\sum_{l\neq i}(\Theta_{l},\mathcal{C}_{l})$
is fully determined by the pair $\bigl(\mathbf{T},\bm{\breve{\mathcal{C}}}\bigr)$, in an explicit way.
Proof.
This follows from Propositions 2.16 and 2.17. Indeed, applying them to determine the invariants of $F\cdot\bigl(\Theta_{i}^{<2},\mathcal{C}_{i}\bigr)$ in terms of those of $\bigl(\Theta_{i}^{<2},\mathcal{C}_{i}\bigr)$, we obtain that
  • $\bullet$
    The fission datum of the global modified formal data $F\cdot\bigl(\Theta_{i}^{<2},\mathcal{C}_{i}\bigr)+\sum_{l\neq i}(\Theta_{l},\mathcal{C}_{l})$ is fully determined by $\mathbf{T}$, in an explicit way.
  • $\bullet$
    The slope of any Stokes circle $I_{a}$ at a singularity $a\neq\infty$ at finite distance of $F\cdot\bigl(\Theta_{i}^{<2},\mathcal{C}_{i}\bigr)$ is determined by $\mathbf{T}$ in an explicit way. In turn, this determines the fission exponent $f_{I_{a},\langle 0\rangle_{a}}$ between $I_{a}$ and the tame circle at $a$, and thus determines the fission datum of the nonmodified formal data associated to $F\cdot\bigl(\Theta_{i}^{<2},\mathcal{C}_{i}\bigr)+\sum_{l\neq i}(\Theta_{l},\mathcal{C}_{l})$.
This makes it possible to take this as a definition of the classes of nearby representations of the enriched tree $\bigl(\mathbf{T},\bm{\breve{\mathcal{C}}}\bigr)$.
@@
Let $\bigl(\mathbf{T},\bm{\breve{\mathcal{C}}}\bigr)$ be an enriched tree, i.e., $\mathbf{T}$ is a short fission tree, $N$ its set of leaves and $\bm{\breve{\mathcal{C}}}$ the datum of a conjugacy class $\mathcal{C}_{i}\in\operatorname{GL}_{n_{i}}(\mathbb{C})$ for each leaf $i\in N$, where $n_{i}$ is the multiplicity of $i$. Assume that there exists an irregular class $\Theta_{\infty}$ of generic form with fission tree $\mathbf{T}$ such that $\bigl(\Theta_{\infty},\bm{\breve{\mathcal{C}}}\bigr)$ are effective modified global formal data. For $i\in\{1,\dots,k\}$, keeping previous notations, we define the $i$-th nongeneric class of nearby representations of $\bigl(\mathbf{T},\bm{\breve{\mathcal{C}}}\bigr)$ as the class $(\mathbf{F}_{i},\bm{\mathcal{C}}_{i})$ of admissible deformations of the nonmodified formal data associated to the formal data
$F\cdot\bigl(\Theta_{i}^{<2},\mathcal{C}_{i}\bigr)+\sum_{l\neq i}(\Theta_{l},\mathcal{C}_{l}).$
(This is well-defined since by the previous proposition this does not depend on the choice of $\Theta_{\infty}$.)
Theorem 3.8 can then be reformulated in the following way, which justifies our terminology.
@@
Let $(\mathbf{\Theta},\bm{\mathcal{C}})$ be an effective wild Riemann surface with boundary data on $\mathbb{P}^{1}$ of generic form, with class of admissible deformations $(\{\mathbf{T}\},\bm{\mathcal{C}})$, and let $(\mathbf{F}_{i},\bm{\mathcal{C}_{i}})$, with $i\in\{1,\dots,k\}$, be the classes of nearby representations of $(\mathbf{T},\bm{\mathcal{C}})$ (in the sense of the previous definition). Then the classes of representations of the nonabelian Hodge space $\mathcal{M}(\mathbf{\Theta},\bm{\mathcal{C}})$ which are nearby $(\mathbf{\Theta},\bm{\mathcal{C}})$ (in the sense of Section 1.3) are exactly $(\{\mathbf{T}\},\bm{\mathcal{C}})$ and $(\mathbf{F}_{i},\bm{\mathcal{C}}_{i})$ for $i\in\{1,\dots,k\}$.
By abuse of language, we will sometimes allow ourselves to say that $\{\mathbf{T}\}$ and $\mathbf{F}_{i}$, for $i\in\{1,\dots,k\}$ are the nearby representations of $\mathbf{T}$.
@@
Let $(E,\nabla)$ be an irreducible algebraic connection on a Zariski open subset of $\mathbb{P}^{1}$, such that $(E,\nabla)$ is not a rank one connection with a unique singularity at infinity with Poincaré–Katz rank $\leq 1$. Let $\bigl(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}}\bigr)$ be its modified global formal data, and $A\in\operatorname{SL}_{2}(\mathbb{C})$ such that $A\cdot\bigl(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}}\bigr)$ is of generic form. Let $\mathbf{T}$ be the ((short)) fission tree of $A\cdot\bm{\breve{\Theta}}$, and $\bm{\mathcal{C}}$ the collection of conjugacy classes of $A\cdot\bigl(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}}\bigr)$. Then the pair $(\mathbf{T},\bm{\mathcal{C}})$ does not depend on the choice of $A$.
@@
This is an immediate consequence of Theorem 3.8. ∎
@@
Keeping the setup of the proposition, we say that $(\mathbf{T},\bm{\mathcal{C}})$ is the enriched tree of $(E,\nabla)$ and write $\mathscr{T}(E,\nabla):=(\mathbf{T},\bm{\mathcal{C}})$.
Proposition 3.12 implies that the enriched tree is invariant under symplectic transformations.
@@
If $(E,\nabla)$ is irreducible and is not a rank one connection with a unique singularity at infinity of Poincaré–Katz rank $\leq 1$, so that $A\cdot(E,\nabla)$ is well-defined for any $A\in\operatorname{SL}_{2}(\mathbb{C})$, we have $\mathscr{T}(A\cdot(E,\nabla))=\mathscr{T}(E,\nabla)$.
3.3 Properties of classes of nearby representations
Proposition 3.9 gives an explicit method to determine the nearby classes of representations from an enriched tree, using Propositions 2.16 and 2.17. We will compute many examples in Section 4. Before this, in this paragraph, let us state several general properties satisfied by the classes of nearby representations.
First, we can get explicit formulas for the rank of the nearby representations. For $i\in\{1,\dots,k\}$, let us denote by $\bigl\langle q^{+}_{i,j}\bigr\rangle$ the elements of $N_{i}^{+}$, and write $\bigl\langle q^{+}_{i,j}\bigr\rangle=\bigl\langle-\frac{a_{i}}{2}z^{2}+q^{\prime+}_{ij}\bigr\rangle$, with $q^{\prime+}_{ij}$ having ramification $r^{+}_{ij}$ and slope $s^{+}_{ij}/r^{+}_{ij}>1$. Similarly, let us denote by $\bigl\langle q^{-}_{i,j}\bigr\rangle$ the elements of $N_{i}^{-}$, and write $\bigl\langle q^{-}_{i,j}\bigr\rangle=\bigl\langle-\frac{a_{i}}{2}z^{2}+b_{ij}z+q^{\prime-}_{ij}\bigr\rangle$, with $q^{\prime-}_{ij}$ having ramification $r^{-}_{ij}$ and slope $s^{-}_{ij}/r^{-}_{ij}<1$. Let $n_{ij}^{\pm}$ be the multiplicities of the active circles.